Write and equation for each sentence. Is each equation true, false, or an open sentence?
Thirty-two plus five equals the product of six and six.
A number t divided by seventeen equals the opposite of three.
The product of four and twenty equals eighty.

Answer by stanbon(75887)

Write and equation for each sentence. Is each equation true, false, or an open sentence?
Thirty-two plus five equals the product of six and six.
32+5 = 6*6
False
---------------
A number t divided by seventeen equals the opposite of three.
t/16 = -3
Open
--------------------
The product of four and twenty equals eighty.
4*20 = 80
True
=============
Cheers,
Stan H.